OpenStudy (ammarah):

Do u know how to convert 380.0 degrees to radian measure?

OpenStudy (ammarah):

@ganeshie8

OpenStudy (ammarah):

@whpalmer4

OpenStudy (displayerror):

Multiply the angle (in degrees) by pi/180 \[380 \deg \times \frac{\pi}{180 \deg}\]